# All-Optical Brillouin Random number Generator

### v1

@@ V Optomechanical Random Number Generator
To set up the device, firstly we estimate the pump amplitude of $\Omega_{1}$ where average lifetimes of the generating and the non-generating states ($\tau_{g}$ and $\tau_{ng}$) are equal to each other. Secondly, we estimate a boundary amplitude value of the second mode, $|a_{2bound}|$ [9, 10], which separated the generating and the non-generating states. When the amplitude of the second optical mode is greater than the boundary value ($|a_{2}|>|a_{2bound}|$), it means that the current state is the generating one and formally can be considered as 11. In opposite case when $|a_{2}|<|a_{2bound}|$, the current state is the non-generating one which can be considered as 0. Finally, we estimate the sampling frequency. In the case when the average lifetime of the two states are equal ($\tau_{g}=\tau_{ng}=\tau$) we can introduce a half-life time $T_{1/2}=\tau\log{2}$. Here the half-life time $T_{1/2}$ is the lifetime of the both states: the non-generating and the generating ones. NIST tests required a balanced bit stream with equal amount of 0 and 11. The balance of bit stream can be setup by the amount of $|a_{2bound}|$. Furthermore, some NIST tests to prove randomly generated data required frequent transitions between states. We can control the transition frequency by the amplitude of the external seed wave, $\Omega_{2}$, or by varying the sampling frequency. In our model, to observe more frequent transitions between 0, 11 and pass all NIST test, we use the next sampling frequency $f_{s}=\dfrac{1}{4T_{1/2}}$. At this sampling frequency, the generator passes all NIST tests (see Appendix). Using parameters from the Fig. 2, we obtain that the speed of bit generating is about 55 Mb/s.
@@ VI Conclusion
In summary, we have investigated the Brillouin optomechanical system operating in the hard excitation mode as an all-optical device for true random number generation. Using thermal noise and spontaneous transitions between two stable states, we have shown that in such system the probability of being in the states and the average lifetimes of the states can be controlled by the external pump and seed waves. The explanation of such phenomena because of the relative stability of the generating and non-generating states depend on the external waves. This allows to work in the regime where the probabilities of being in the states are equal to each other ($p_{ng}=p_{g}=0.5$). This device can be used as random number generator without any digital data processing and have possible realization on chip. It provides a random bit generation which passes all standard NIST tests with the frequency of generation of about 55 Mb/s.

### v2

@@ V Optomechanical Random Number Generator
To set up the device, firstly we estimate the pump amplitude of $\Omega_{1}$ where average lifetimes of the generating and the non-generating states ($\tau_{g}$ and $\tau_{ng}$) are equal to each other. Secondly, we estimate a boundary amplitude value of the second mode, $|a_{2bound}|$ [9, 10], which separated the generating and the non-generating states. When the amplitude of the second optical mode is greater than the boundary value ($|a_{2}|>|a_{2bound}|$), it means that the current state is the generating one and formally can be considered as $1$. In opposite case when $|a_{2}|<|a_{2bound}|$, the current state is the non-generating one which can be considered as 0. Finally, we estimate the sampling frequency. In the case when the average lifetime of the two states are equal ($\tau_{g}=\tau_{ng}=\tau$) we can introduce a half-life time $T_{1/2}=\tau\log{2}$. Here the half-life time $T_{1/2}$ is the lifetime of the both states: the non-generating and the generating ones. NIST tests required a balanced bit stream with equal amount of 0 and $1$. The balance of bit stream can be setup by the amount of $|a_{2bound}|$. Furthermore, some NIST tests to prove randomly generated data required frequent transitions between states. We can control the transition frequency by the amplitude of the external seed wave, $\Omega_{2}$, or by varying the sampling frequency. In our model, to observe more frequent transitions between 0, 11 and pass all NIST test, we use the next sampling frequency $f_{s}=\dfrac{1}{4T_{1/2}}$. At this sampling frequency, the generator passes all NIST tests (see Appendix). Using parameters from the Fig. 2, we obtain that the speed of bit generating is about $5$ Mb/s.
@@ VI Conclusion
In summary, we have investigated the Brillouin optomechanical system operating in the hard excitation mode as an all-optical device for true random number generation. Using thermal noise and spontaneous transitions between two stable states, we have shown that in such system the probability of being in the states and the average lifetimes of the states can be controlled by the external pump and seed waves. The explanation of such phenomena because of the relative stability of the generating and non-generating states depend on the external waves. This allows to work in the regime where the probabilities of being in the states are equal to each other ($p_{ng}=p_{g}=0.5$). This device can be used as random number generator without any digital data processing and have possible realization on chip. It provides a random bit generation which passes all standard NIST tests with the frequency of generation of about $5$ Mb/s.